# Token Preference Optimization with Self-Calibrated Visual-Anchored Rewards for Hallucination Mitigation

### v3

@@ A.3 Training Efficiency
It has also shown superior outcomes on 5K training data training to CSR training on 13K data and POVID training on 17K data. This highlights the efficacy of our method in guiding the model to pay more attention to image details and in reducing hallucinations. We promise we will further elaborate on our efficiency in detail in the final version.
Appendix B Ablation Analysis
Noise Step
We ablate on the noise steps in Figure 5 (a). As shown, the optimal performance is achieved at the step of 500. This medium corruption enables the model to grasp the general outline of the image while missing the detailed contents, which is prone to generate hallucinations of the visual-anchored tokens.
Parameter $\bf a$
We conduct experiments by varying the parameter $a$ introduced in Equation (5) with the results shown in Figure 5 (b). By setting $a=[0,0.5,1]$, we observed consistently good performance across all configurations. This suggests that effective performance is achieved as long as the reward mechanism successfully highlights token differences and identifies visually anchored tokens. Notably, the best overall results are obtained with $a=0.5$, validating our proposed method and hypothesis. This indicates that when the visual-anchored score $s=0$, setting $c=1$, not introducing additional reward signals can yield better outcomes.
@@ Appendix B Ablation Analysis
Appendix C Comparison with Decoding-based Methods
In addition to preference alignment, many researches have improved the decoding strategy of the model to eliminate hallucinations. In this section, we selected some typical works on decoding strategies, such as VCD Leng et al. (2024), OPERA Huang et al. (2024) and LURE Zhou et al. (2023), and compared the results with them. The results show that TPO achieves more significant hallucination reduction.
@@ Appendix C Comparison with Decoding-based Methods
Appendix D Results on Object Hallucination
@@
The results are shown in Figure 6. Note that ’Chair’ represents the hallucination ratio, where a smaller value indicates better model performance. To more clearly illustrate the comparison between methods in the figure, we use $10-chair$ as the indicator. The results show that TPO can not only mitigate the hallucination in visual question answering, but also eliminate the object hallucination in image descriptions to a certain extent.
@@
Appendix E Comparison of Different Noise Adding Methods.

### v4

@@ A.3 Training Efficiency
It has also shown superior outcomes on 5K training data training to CSR training on 13K data and POVID training on 17K data. This highlights the efficacy of our method in guiding the model to pay more attention to image details and in reducing hallucinations.
@@
Appendix B Comparison with Related Methods
To more comprehensively highlight the advantages of the TPO method, we conducted comparisons with other related works Jiang et al. (2024); Yue et al. (2024); Xiao et al. (2025); Sarkar et al. (2024); Zhao et al. (2023); Leng et al. (2024); Huang et al. (2024); Zhou et al. (2023) aimed at addressing the hallucination problem. The results show that TPO achieves more significant hallucination reduction.
@@ Appendix B Comparison with Related Methods
To further elucidate the innovations of our work, we provide a more in-depth comparison with similar methods Cui et al. (2024); Zhou et al. (2024a); Leng et al. (2024). Our approach exhibits both conceptual distinctions and methodological advancements. Unlike FiSAO, which relies on a vision encoder to compute token-level scores, TPO derives them directly from the LVLM’s logits difference between clean and noisy images. Moreover, while FiSAO builds upon PPO, TPO enhances the DPO framework. Our method shares with VCD the idea of using noise-induced logits change to identify visual-sensitive tokens; however, TPO goes beyond identification by introducing a self-calibrating reward Eq. 5 that jointly reinforces positive outputs, suppresses hallucinations, and strengthens image–token alignment—aiming fundamentally to improve perceptual faithfulness. In contrast, VCD focuses on decoding less image-affected tokens without enhancing underlying visual grounding. Similarly, although POVID also uses Gaussian noise for perturbation, it primarily constructs negative samples for DPO training, whereas TPO leverages noise to quantify token–image relevance and enables a more efficient token-level reinforcement learning strategy, leading to superior performance.
Appendix C Results on Object Hallucination
@@
The results are shown in Figure 6. Note that ’Chair’ represents the hallucination ratio, where a smaller value indicates better model performance. To more clearly illustrate the comparison between methods in the figure, we use 10chair10-chair as the indicator. The results show that TPO can not only mitigate the hallucination in visual question answering, but also eliminate the object hallucination in image descriptions to a certain extent.
@@
Appendix D Comparison of Different Noise Adding Methods.
@@
Appendix E Evaluation of TPO Training on Real-World Performance
To validate the effect of TPO training, we compared the performance of the LLaVA-1.5-7B models before and after training on a real-world image set of 10 images, as suggested during the review process. A manual analysis of the outputs revealed that the TPO-trained model provided superior results in 40% of the cases. In contrast, the original model was better in only 20% of cases, with both models performing equally well in the remaining 40%. This represents a clear net performance gain attributable to the TPO training procedure.
Method AMBER
Acc F1
LLaVA-1.5-7B 71.7 74.3
+ DPO 77.5 82.1
+ CSR 73.2 76.1
+ POVID 71.9 74.7
+ RLHF-V 74.8 78.5
+ TPO (1K) 72.5 75.3
+ TPO (3K) 78.9 83.4
+ TPO (5K) 79.3 85.0
Table 9: Evaluation of TPO Training with varying amounts of training data.
Appendix F Analysis of TPO Training on less training data
In order to further investigate the impact of training data volume, we constructed subsets of 1K and 3K samples from the original 5K training set. Using these subsets, we fine-tuned the LLaVA-1.5-7B model with a consistent learning rate of 5e-8 for 4 epochs, maintaining identical experimental settings as in the full-data scenario.
As anticipated in table 9, the performance of the TPO method exhibits a gradual decline with reduced training data. Nevertheless, even when trained on only 1K samples, TPO remains competitive and surpasses most baseline methods. With 3K training samples, it achieves the second-highest performance, further underscoring the effectiveness and data efficiency of our approach.